OpenStudy (anonymous):

13=-7+x X= 20 is this conditioonal contradiction identiity show me how you got your steps?? im trying to learn process

OpenStudy (anonymous):

13 = -7 + x 13(+7) = -7(+7) + x 20 = 0+x 20=x HTH